OpenStudy (anonymous):

what is the area of a square with sides of length 6?

OpenStudy (anonymous):

6*6=36

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

it's not 24?

OpenStudy (mathstudent55):

For a square, A = s^2, where A = area, and s = length of side

OpenStudy (mathstudent55):

A = 6^2 = 6 * 6 = 36

OpenStudy (anonymous):

no, it's not 24 :)

OpenStudy (anonymous):

Oh but is an area of a square P=4s?

jimthompson5910 (jim_thompson5910):

P = 4s is the perimeter, not the area, of a square

OpenStudy (anonymous):

it's not area. it's perimeter :)

OpenStudy (anonymous):

OHH that's right hahah

OpenStudy (anonymous):

that you guys!!

OpenStudy (skullpatrol):

Not funny.

OpenStudy (anonymous):

but you're very funny @skullpatrol :)

OpenStudy (anonymous):

who asked you @skullpatrol ?

jimthompson5910 (jim_thompson5910):

i'm sure it's harmless skullpatrol, stop being a "square" lol

jimthompson5910 (jim_thompson5910):

sry lame math/nerd humor...

OpenStudy (anonymous):

@jim_thompson5910 @marsss lol xD it's all cool!

OpenStudy (skullpatrol):

What about me :'(

OpenStudy (anonymous):

@skullpatrol You're awesome too :)

OpenStudy (anonymous):

he's just a fool :)

OpenStudy (anonymous):

@marsss awh lol thank you guys for helping me out on this, really! Without ya'll, I don't know if I could do it!!

OpenStudy (anonymous):

welcome, princess :)